Find the inverse of 55 modulo 7 by using extended Euclidean Algorithm



step by step solution

1
Expert's answer
2021-11-08T12:07:23-0500

Apply Euclidean Algorithm to compute GCD as shown in the left column.

It will verify that "GCD(55, 7)=1."

Then we will solve for the remainders in the right column.


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 55=7(7)+6 & 6=55-7(7) \\ 
 7=6(1)+1 & 1=7-6(1) \\
 6=1(6)+0 & 
\end{array}"

Use the equations in the right side and perform reverse operation as:


"\begin{matrix}
 1=7-6(1)\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \\ \\
 \ \ \ 1=7-[55-7(7)](1)\\ \\
1=7(8)+55(-1)\ \ \ 
\end{matrix}"

Therefore "1=55(-1)\mod {7}," or of we prefer a residue value for multiplicative inverse


"1=55(6)\mod {7}."

Therefore, "7" is the multiplicative inverse of "55."